OpenStudy (anonymous):

Michael is helping his mom make decorations for his grandparent's 50th wedding anniversary celebration. They've decided to make paper chain decorations to put around the room. So far, Michael has used 15 paper links to make a chain 3 feet long. How many paper links will he need to make 4 chains that are each 10 feet long?

OpenStudy (hesan):

Since 3 feet chain made by 15 papers Therefore 1 foot chain made by 15/3 = 5 papers and 10 feet chain will be of 5 * 10 = 50 papers 4 chains of 10 feet each uses 50 * 4 = 200 papers. SO the answer is 200 papers..